OpenStudy (anonymous):

use the quadratic formula to find any x-intercepts on the graph of the equation. y=5x^2+6x-6

OpenStudy (mertsj):

Found it.

OpenStudy (mertsj):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Thank you!!!!!!!!!

OpenStudy (mertsj):

That is the quadratic formula. In your problem a =5, b = 6, c = -6 Plug in those values and solve.

OpenStudy (anonymous):

i wish I post what I got so far but I dont have the format

OpenStudy (mertsj):

Use the equation button below this box

OpenStudy (anonymous):

DUH!!!

OpenStudy (anonymous):

\[\frac{ -6\pm \sqrt{45} }{ 10 }\]

OpenStudy (mertsj):

\[x=\frac{-6\pm \sqrt{6^2-4(5)(-6)}}{2(5)}=\frac{-6\pm \sqrt{36+120}}{10}\]

OpenStudy (anonymous):

how?!

OpenStudy (mertsj):

Negative times negative = positive

OpenStudy (anonymous):

oooooh!! I messed up

OpenStudy (anonymous):

\[\frac{ -6\pm \sqrt{156} }{ 10 }\]